The $/€ bid and ask prices are $1.45/€ and $1.48/€, respectively. The €/$ bid and ask prices are:

Multiple Choice

  • $1.48/€1 and $1.45/€1.

  • €0.6897/$1 and €0.6757/$1

  • €0.6757/$1 and €0.6897/$1.

  • €0.6623/$1 and €0.6667/$1.

  • None of the options.
Answer #1

Euro/Dollar Bid price = 1/1.48 = Euro 0.6757/$1

Ask price = 1/1.45 = Euro 0.6897/$1

Hence, the answer is

  • €0.6757/$1 and €0.6897/$1